0.1 Frequently Asked Questions

This is the Emacs/W3 Frequently Asked Questions list.

0.1.1 Installation

• Q1.1 What is the latest version of Emacs/W3?

The latest version is v4.0pre.46.

• Q1.2 Where do I get Emacs/W3?

Any of the following locations:

 − http://www.cs.indiana.edu/elisp/w3/download.html
 − ftp://ftp.cs.indiana.edu/pub/elisp/w3/

• Q1.3 Which version of Emacs do I need?

At least GNU Emacs 19.34, or XEmacs 19.14 is recommended. GNU Emacs 19.28 has been known to work, but highlighting will not work due to differences in how faces are handled in newer Emacs versions.

• Q1.4 I get ’./configure: sh internal 2K buffer overflow’ when configuring. On some systems (notably HP-UX 9.x), the standard /bin/sh is not quite up to snuff. You should use bash to run configure if you possibly can, ie: bash ./configure [standardarguments] If anyone has ideas about how I could fix my autoconf sources, please contact

mailto:wmperry@aventail.com.

• Q1.5 How can I build Emacs/W3 under Windows 95 or Windows NT?

• Q1.6 How can I build Emacs/W3 under VMS?

• Q1.7 I still want to use Emacs 19, are there any special problems? Since Emacs 19.x does not include a suitably ‘new’ version of the custom and widget libraries, you will need to install them yourself. Please see

http://www.dina.kvl.dk/~abraham/custom/ for more information on in-

stalling. After this, you will need to re-run your configure script, including the –with-custom=/path/to/custom/library/directory.

0.1.2 Runtime

• Q2.1 The options menu is ugly under Emacs 19 In Emacs 19, Emacs/W3 uses the ’easymenu’ package to define menus. Unfortunately, this package does not show the current state of a boolean menu entry, it just shows as ’Toggle xxxx’. This feature-lack will be fixed in Emacs 19.35 or Emacs 20.1 when they are released.

• Q2.2 Animated GIFs cause XEmacs 19.14 or 20.0 to crash.

There was a bug in the GIF decoding routines in XEmacs 19.14 and 20.0 - please update to XEmacs 19.15 or 20.1 respectively.

• Q2.3 I get an error when starting up about custom-declare-group This usually means you compiled Emacs/W3 in an emacs that could not find the ’custom’ package (or found an old version), but you are running it in an emacs that finds the new one. This is usually the case if you have a recent version of GNUS installed in a non-standard place. Make sure that you can load the custom library when compiling emacs. Set the environment variable WIDGETDIR to where your custom library lives (ie: ~/lisp/gnus/lisp), or rerun configure with the ’–with-custom=/full/path/to/custom/lisp/’ command line argument and then remake with a ’make clean all’.

• Q2.4 It is incredibly slow under OS/2 This appears to be a problem with how fonts are looked up. Thanks to chang@wsu.edu for this analysis. In OS/2, the font name default to forms like

     "-*-Courier-medium-r-normal--*-100-*-*-m-*-cp850" ......
     or
     "10.Courier"

In these two cases, font detection fails because this does not match x-font-regexp. To solve this, set the default font as "-*-Courier-medium-r-normal–*-100-*-*-m-*- cp850-1"

• Q2.5 How do I make emacs scroll the window horizontally when moving?

In XEmacs, you should simply put something like this in your .emacs file:

        (add-hook 'w3-mode-hook '(lambda () (auto-show-mode 1)))

Under Emacs 19, there are a variety of different packages that provide this functionality. The one I have seen used the most is ’hscroll.el’, from

http://reality.sgi.com/wmesard/ftp/emacs/hscroll.el - to turn it on, add this

to your .emacs file:

        (autoload 'turn-on-hscroll "hscroll" nil t)
        (add-hook 'w3-mode-hook 'turn-on-hscroll)

• Q2.6 I like being warned about invalid HTML on my own pages, but how can I make Emacs/W3 stop telling me almost everything on the Web is invalid? Snide answer: Make everyone start writing valid HTML pages. :) Real answer: You can use the file preparation hook, which is run before any parsing is done. Something like this should work.

     (defun my-w3-file-prepare-hook ()
       (make-local-variable 'w3-debug-html)
       (setq w3-debug-html
           (if (or (string= (url-type url-current-object) "file")
                   (string-match ".*\\.some\\.domain\\.name"
                                 (or (url-host url-current-object) "")))
               'style
             nil)))
     (add-hook 'w3-file-prepare-hook 'my-w3-file-prepare-hook)

This will turn on stylistic warnings for any local HTML files or files loaded from the ‘*.some.domain.name’ domain.

• Q2.7 I get an error restarting Emacs after using the preferences menu in Emacs/W3:

‘Symbol's function definition is void: custom-set-variables’

Typically this happens if you are running GNU Emacs 19. Because the custom libraries are not bundled with Emacs, you need to make sure that you load it from within your .emacs file. Place the following line before any of the calls to custom-set-variables:

     (require 'custom)

• Q2.8 When using Emacs 20.x, a \201 character appears in the buffer before every character above ASCII 127? You should install the LEIM package and put the following lines in your .emacs file:

     (set-language-environment 'Latin-1)
     (standard-display-european t 'iso-latin-1)

LEIM packages for Emacs 20.x are located at ftp://prep.ai.mit.edu/pub/gnu/, filename is ‘leim-XX.Y.tar.gz’, where XX.Y matches the Emacs version you are currently running.

• Q2.9 All my FTP directory listings show up as HTML! I HATE THAT!

This will happen if you are going through a web-based proxy server. To use the normal ftp support that Emacs provides, you will need to set up the no-proxy section of your proxy services to not include any ftp traffic See section “Proxy Gateways” in Emacs/W3 User’s Manual. You must then also configure Ange-FTP or EFS to know about your specific proxy setup. After all this pain and anguish, you will see your FTP directories in dired as the gods intended.

• Q2.10 Why can’t I choose an option in an HTML form using the minibuffer instead of the #%!@ing ’widget choose’ buffer? This is in the hands of the widget library. Thankfully, there is a way to change this

behaviour. Set the variable widget-menu-minibuffer-flag to t in your ‘~/.emacs’

file, and you will get the expected behaviour.

0.1.3 Arcana

• Q3.2 How do i get Shift-Tab to go backwards on a text terminal or XTerm? aka: I hate the new text widgets, I can’t go through the links with n and b I can go forward using TAB but how do i go backward on a terminal? Thanks to Greg Stark <gsstark@mit.edu> for this incredibly thorough answer. I have verified that this works on my Linux laptop. Not all terminals can distinguish between a shifted tab and an unshifted tab at all. Tab is indicated on a text terminal by a control-i. There is no such thing as capital control characters, so if the terminal is going to indicate a shift-tab somehow it has to be completely differently. The most appropriate thing to use is probably "backtab" which on old text terminals was sometimes a separate key and Emacs is already set up to recognize automatically if it exists. Making "backtab" work involves several steps. First you have to make sure your console generates some character sequence to indicate the key you want to generate a "backtab". Then you have to configure termcap or terminfo to recognize that key sequence. Then you may have to make your programs do useful things when they get a "backtab", Emacs for example will recognize it automatically but except for the Widget and W3 commands nothing is ever bound it it.

Step 1 On An XTerm: XTerm obeys standard X Toolkit translations which you can use to specify what character sequence Shift-Tab generates. The following X Resources will cause Shift-Tab and Meta-Shift-Tab to generate reasonable character sequences. You can either put this in your .Xresources or .Xdefaults file, or you can put it in /usr/lib/X11/app-defaults/XTerm to make it a site-wide default. (On Debian systems you should put it in /etc/X11/Xresources, not the app-defaults files):

     XTerm*VT100.translations: #override \
        ~Meta Shift<Key>Tab: string(\033[Z) \n\
         Meta Shift<Key>Tab: string(\033\033[Z) \n

I recommend these sequences, they are based on what seems to be a more or less standard sequence ^[[Z for backtab. Step 1 On Rxvt: By default Rxvt sends ^[[Z for Shift-Tab. However, if Shift-Tab generates another keysym, like for example in XFree86 3.2 where it’s bound to ISO Left Tab then Rxvt will just ignore it. You would need to defeat this feature to make rxvt work again by doing something like:

     xmodmap -e 'keysym Tab = Tab'

or adding that command to some global X configuration file (On Debian systems adding “Keysym Tab = Tab” to /etc/X11/Xmodmap or ~/.Xmodmap is sufficient) Step 1 On A Linux Virtual Console: on a Linux virtual console you can configure what character sequences are generated by which keys using the loadkeys command. Many systems are set up to run loadkeys automatically on startup with some keymap file. On Debian systems this is true, the keymap file is specified in /etc/kbd/config and usually lives in the /usr/lib/kbd/keytables directory. You want to put something like the following in your keytable file:

     keycode  15 = Tab   F91
     alt     keycode  15 = Meta_Tab
     shift   alt     keycode  15 = F92

where keycode 15 is Tab on my keyboard (and probably any keyboard). This defines Tab and Alt-Tab normally, and also defines Shift-Tab to be F91 and Shift-Alt-Tab to be F92. Then put something like this:

     # backtab and M-backtab
     string F91 = "\033[Z"
     string F92 = "\033\033[Z"

later in the file. This defines what character sequence F91 (Shift-Tab) and F92 (Alt-Shift-Tab) should generate. I recommend these sequences, they are based on what seems to be a more or less standard sequence ^[[Z for backtab. Step 2 On A Termcap System: The termcap capability is kB, i’m not familiar with termcap tools, i think you just need to add it to the /etc/termcap file for the terminal you’re concerned with as kB=\E[Z. Step 2 On A Terminfo System: The terminfo capability is kcbt (the long name is key btab). You want to run infocmp to generate an edittable copy of the terminal info. Add the capability, then use tic to compile that information. Something like this: infocmp $TERM > info emacs info & # add kcbt=\E[Z, to the file tic info If you do this as root it should add the new definition to the system wide terminfo database. If you do it as a normal user it should create a ~/.terminfo database with a local terminfo info definition for that terminal. Step 3 On Emacs: The standard terminal initialization should recognize the backtab capability automatically. To test it try C-h c Shift-Tab and see what it calls the key. To bind commands to it just use [backtab] in local-set-key or global-set-key as in:

     (local-set-key [backtab] 'hippie-expand)
     or
     (global-set-key [backtab] 'hippie-expand)

In the interest of maintaining a single consistent set of key bindings between X and tty emacsen you may want to make equivalent X keystroke generate "backtab" as well, you can do this by doing this: (define-key function-key-map [S-tab] [backtab]) or (define-key function-key-map [isolefttab] [backtab]) To make S-tab or whatever keystroke you made generate backtab on a terminal be recognized as backtab under X11 as well. You can check how Emacs recognizes this keystroke currently by doing C-h c <keystroke>.

0.1.4 Customization